OpenStudy (anonymous):

If the area of the triangle is 36cm2, what is the length, in cm, of leg xy? answer is 12? can someone show me how to work this problem?lost

OpenStudy (anonymous):

what is the formula for the area of a triangle?

OpenStudy (anonymous):

Thats the info I have.

OpenStudy (anonymous):

Is there a diagram showing the triangle? We have to know something else... there is no way to answer the length of xy without a diagram or more info. Area of any triangle = 1/2 times length of the base times height or often, A = (1/2)bh so xy is probably one of the legs, either the base or the height, and they tell you area is equal to 12. but without knowing the other leg, there isn't anything you can do.

OpenStudy (anonymous):

Its a right triangles Leg xy of the right triangle show in the diagrams twice as long as the legX=YZ..IF the area of the triangle is 36cm 2, what is the lenght, in cm, of the leg xy? answer is 12..Show me how to work this problem?

OpenStudy (anonymous):

yes, so you do have more info... by knowing that the base xy is twice the length of the height, yz, you can set up the problem: Area = 36 = (1/2)base x height = (1/2)(length of xy)(length of yz) but since you know that length of xy is twice length of yz, that means the length of yz is half of the length of xy. so let's quit using yz and instead substitute in... A = 36 = (1/2)(length of xy)(1/2 the length of xy) So now, if you reorganize, you have (1/2) times (1/2) times (len of xy) times (len of xy), all equal to 36. Or you have 36 = (1/4)(length of xy)^2 you can multiply both sides by 4, leaving 144 = (length of xy)^2

OpenStudy (anonymous):

What do you get for the length of xy?

OpenStudy (anonymous):

xy is 12, as you know, and that works, since 12*12 = 144 You can also check the area: (1/2)(12)(half of 12) = (1/2)(12)(6) = (1/2)(72) = 36, the area that the problem gave you.

OpenStudy (anonymous):

thanks!:)

OpenStudy (anonymous):

glad to help... hope it made sense after seeing how to work it.

OpenStudy (anonymous):

yes:) thanks